The abiscissa of a point in the 4th quadrant is numerically three times its ordinate and is 10 units from (-2,4). Find the point.


1
Expert's answer
2021-09-23T16:54:04-0400

 Let the required point be "(x,y)."  Then

"x<0, y<0, x=3y"


"d^2= (x-(-2)^2+(y-4)^2=10^2"



"(3y+2)^2+(y-4)^2=100"

"9y^2+12y+4+y^2-8y+16-100=0"

"10y^2+4y-80=0"

"5y^2+2y-40=0"


"D=(2)^2-4(5)(-40)=804"

"y=\dfrac{-2\pm\sqrt{804}}{2(5)}=\dfrac{-1\pm\sqrt{201}}{5}"

Since "y<0," we take "y=\dfrac{-1-\sqrt{201}}{5}."

"x=3(\dfrac{-1-\sqrt{201}}{5})"


Point "(-\dfrac{3+3\sqrt{201}}{5}, -\dfrac{1+\sqrt{201}}{5})"